OpenStudy (anonymous):

7^(x – 2) = 12^x x ≈ 3.28 x ≈ 0.31 x ≈ –7.22 x ≈ –0.14

OpenStudy (mandre):

Just substitute each number and see which one works.

OpenStudy (campbell_st):

you can use logs... base e or base 10 doesn't matter and apply the log laws for powers \[\log(a^b) = b \log a\] applying the log law above \[(x-2)\ln(7) = xln(12)\] distribute the left hand side \[xln(7) - 2\ln(7) = xln(12)\] collect like terms \[-2 \ln(7) = x \ln(12) -x \ln(7)\] factor the right hand side \[-2 \ln(7) = x(\ln(12) - \ln(7))\] apply the log law for division \[\log(\frac{a}{b}) = \log(a) - \log(b)\] so \[-2 \ln(7) = x(\ln(\frac{12}{7})\] hopefully you can solve for x now...